OpenStudy (fernandez555):

Simplify the given expression. (1 + 2i) • (5 − 3i) A. 2 − 15i B. 3 + 2i C.5 − 6i D.11 + 7i

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

i think this is called "multiply" not "simplify"

OpenStudy (anonymous):

\[(1 + 2i) \times (5 − 3i) \] you can either multiply out as you see it, or use \[(a+bi)(c+di)=(ac-bc)+(ad+bc)i\]

OpenStudy (anonymous):

it is in fact D, although getting a letter for an answer from someone here is not going to help you much if you actually have to learn how to do it

OpenStudy (anonymous):

Oh, i just did it really quick